OpenStudy (loser66):

If f:S--> T is a function from S into T, show that the following are equivalent a) f is one-to-one b) For every y in T, the set \(f^{-1({y})\) contains at most one point. c) \(f(D1\cap D2)= f(D1)\cap f(D2)\) for all subsets D1, D2 of S Please, help

OpenStudy (loser66):

@perl

OpenStudy (loser66):

@IMStuck

OpenStudy (anonymous):

is it all equivalent or choose 2 from the 3 ?

OpenStudy (loser66):

each is equivalent to other two

OpenStudy (perl):

so prove a -> b , b - > c , and c - > a that will show statements are equivalent

OpenStudy (loser66):

It shows me nothing @perl

OpenStudy (xapproachesinfinity):

Did you got the question you posted yesterday solved? i looked similar to this no?

OpenStudy (xapproachesinfinity):

it*

OpenStudy (loser66):

No, this is the other one. I didn't solve the yesterday problem yet. I have a bunch of homework to to. I just need a hint or an outline on the problems I think I get stuck. I will go over all of them later.

OpenStudy (xapproachesinfinity):

I'm not good at this but i think to show that \(f(D1~ intersect ~ D2)=f(D1)~intersect~f(D2)\) we must show that an element y belongs to both

OpenStudy (loser66):

But we don't have to prove it, all we need is proving if f is one to one (1) , then (3) not prove (3) itself.

OpenStudy (xapproachesinfinity):

how would you prove that it is one to one? kind of difficult i see that proving c is much better to start with?

OpenStudy (loser66):

oh, I got what you mean

OpenStudy (loser66):

We don't prove it is one to one. We prove: "if f is one to one, then f(D1 and D2)= f(D1 and f(D2)

OpenStudy (xapproachesinfinity):

and how would justify that? we don't know if f in one to one see we need to work the converse?

OpenStudy (loser66):

Ok, let finish it today (hihihi, I am lazy!!) f(D1 and D2) = {x such that image of x is in (D1 and D2)}

OpenStudy (xapproachesinfinity):

yeah! that's the definition don't know how to proceed further lol. what is this course?

OpenStudy (loser66):

Advanced calculus 1

OpenStudy (loser66):

I still have abstract algebra and probability and statistics reference course. hehehhe.. I am ready to die.

OpenStudy (perl):

hehe

OpenStudy (xapproachesinfinity):

this looks to me more like abstract algebra were are dealing with

OpenStudy (loser66):

@perl, at least, you give me step2, right?

OpenStudy (xapproachesinfinity):

we are*

OpenStudy (perl):

which one are you proving ?

OpenStudy (loser66):

hihihi Actually, the problem is quite obvious, so that the hardest part is how to present it professionally.

OpenStudy (xapproachesinfinity):

what is that part 2? f inverse?

OpenStudy (loser66):

@xapproachesinfinity are you good at abstract?? hhehehe, let me become your fan, I will nag you when I get stuck

OpenStudy (xapproachesinfinity):

no man, not good at it hehe. i'm just sticking my noise in this stuff! i'm just become a calculus man hehe

OpenStudy (loser66):

for part2) we have definition of inverse function which says : f^- = {x in S such that f(x) in T} and f is one to one, which leads us to f(x) is unique for any x in s

OpenStudy (loser66):

that is not one to one @perl

OpenStudy (xapproachesinfinity):

it is still 1 though? why are you taking onto case?

OpenStudy (perl):

i know, im trying to satisfy the hypotheses

OpenStudy (loser66):

That is not onto, either @xapproachesinfinity

OpenStudy (xapproachesinfinity):

oh yeah my bad! i thought 2 was linked

OpenStudy (xapproachesinfinity):

so knowing that f(D1 and D2)=f(D1)andf(D2) doesn't necessarily mean it is one to one as perl showed?

OpenStudy (loser66):

Hey, you guys are testing me?? I am a diligent student. I am not lazy to just ask for answer without knowing anything. hehehehe

OpenStudy (xapproachesinfinity):

dealing with this stuff means you are pretty good hehe

OpenStudy (perl):